OpenStudy (anonymous):

Please help me with my homework, I don't know how to do this. Solve the system using elimination, make sure to show work. 5x + 4y = 12 3x – 3y = 18

OpenStudy (anonymous):

Notice \[ 4y\\ -3y \]

OpenStudy (anonymous):

You want them to cancel. so to do that you would multiply the top by 3 and the bottom by 4.

OpenStudy (anonymous):

that would make them \[ 12y\\ -12y \]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Since \(12y-12y=0\)

OpenStudy (anonymous):

15x +12y = 36 12x -12y = 72 is this right so far

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now add both equations.

OpenStudy (anonymous):

27x = 108 x = 108/27 = 4

OpenStudy (anonymous):

15*4+12y=36 y=-2

OpenStudy (anonymous):

x is 4 and y is -2 right

OpenStudy (anonymous):

The best way to check your work is to plug them into the original equations and see if they're true.

OpenStudy (anonymous):

thanks :D